The length of a rectangle is 8 centimeters less than three times its width. Its area is 35 square centimeters.What is the dimension of the rectangle?
Answer by checkley79(3341)

AREA=LW
L=3W-8
35=(3W-8)W
35=3W^2-8W
3W^2-8W-35=0
(3W+7)(W-5)=0
W-5=0
W=5 ANS FOR THE WIDTH.
L=3*5-8
L=15-8=7 ANS. FOR THE LENGTH.
PROOF:
35=7*5
35=35